OpenStudy (anonymous):

Please help me, I need to find the radius and center of 3x^2+3y^2 -6x +12y=0. I grouped the x's and y's together and then factored, but now I am lost.

OpenStudy (anonymous):

i thin u have to factor a 3 then complete the square

OpenStudy (anonymous):

\[3x ^{2} + 3y ^{2} -6x +12y=0\] \[3x(x-2) + 3y(y+4) = 0\]

OpenStudy (anonymous):

I do not know what to do now.

OpenStudy (phi):

The first step is to divide both sides of the equation by 3 divide each term by 3

OpenStudy (phi):

you do not factor 3x^2 -6x you complete the square

OpenStudy (anonymous):

I got (x-1) + (y+2)=5

OpenStudy (anonymous):

I understand it a little bit now. Thank you.

OpenStudy (phi):

you keep leaving out the exponents. Let's see what we get if we expand your answer (x-1)^2 + (y+2)^2 = 5 x^2 -2x+1 + y^2 +4y +4 -5=0 x^2 -2x +y^2 +4y =0 now if we multiply both sides by 3 we get 3x^2 -6x +3y^2 +12y =0 it matches. So, except for leaving out ^2 your have the correct answer

OpenStudy (anonymous):

what would I do if \[2x ^{2} +2y ^{2}-4x=0 \] was my equation?

OpenStudy (anonymous):

since there are only 3 terms, I cannot do what I just did correct?

OpenStudy (phi):

You would divide both sides of the equation by 2 to get x^2 +y^2 -2x =0 complete the square for the x's the y is already a perfect square (remember, the equation of a circle at (0,0) is x^2 + y^2 = r^2 or if you like (x-0)^2 + (y-0)^2 = r^2 your circle's center will have a y=0, and we have to find the x, by completing the square

OpenStudy (anonymous):

(x-1)+(y-0) =1?

OpenStudy (phi):

I will say correct if you write it this way (x-1)^2+(y-0)^2 =1

OpenStudy (anonymous):

lol (x-1)^2 +(y-0)^2=1 thank you

OpenStudy (phi):

and people don't bother to write (y-0) I put it there so you can "see" that it matches our pattern. so on a test you would see a choice like this: (x-1)^2 + y^2=1

OpenStudy (anonymous):

that's good to know

OpenStudy (anonymous):

can I ask you one more question?

OpenStudy (phi):

so your circle has a center at (1,0) and a radius of 1. We can use wolfram to check the answer http://www.wolframalpha.com/input/?i=2x%5E2%2B2y%5E2 −4x%3D0 click on properties to the right of where it says circle

OpenStudy (phi):

Please make your new question a separate post.

OpenStudy (phi):

The wolfram link needs that extra bit -4x%3D0

OpenStudy (anonymous):

thank you